Problem: at time t = 0, a 30-year bond loan (mortgage) is issued.
Coupon rate: 3%
EAR: 1,7%
Face value: 200000

Calculate the price of the bond, assuming it is an annuity bond, per $100, at the time of issue
Answer #1

N = 30

I/Y = 1.7

PMT = 200,000 * 0.03 = 6,000

FV = 200,000

CPT PV

PV = -260,706.1580

Price of the bond per $100 face value = 260,706.1580 * 100/200,000 = $130.353079